A business manager wanted to know if there was a relationship between daily temperature measured in degrees Fahrenheit and number of employees who called-in sick each day. He collected the following data over 10 days:

Daily Temperature: 64 60 67 63 60 64 72 66 60 69

# Called-in Sick: 33 32 30 37 34 28 31 27 31 26

Use the Daily Temperature (X-variable) and Called-in Sick (Y-Variable) Data to answer the following questions.

a. What is the value of the slope of the regression line?

b. What is the value of the y-intercept?

c. Provide an interpretation of your answer to part B.

d. What is the final form of the regression line (i.e., y’ = a + b(x))?

e. Use Excel to obtain a scatterplot of the data with the regression line embedded in it.

f. What is the predicted number of employees that will call-in sick on a day that is 71 degrees Fahrenheit?
Answer #1
x y (x-\bar{x})(y-\bar{y}) (x-\bar{x})^2
64 33 -3.36 2.56
60 32 -6.16 31.36
67 30 -1.26 1.96
63 37 -15.86 6.76
60 34 -17.36 31.36
64 28 4.64 2.56
72 31 0.64 40.96
77 27 -44.46 129.96
60 31 -0.56 31.36
69 26 -16.66 11.56
Total 656 309 -100.4 290.4
Total/n 65.6 30.9

See that the last column, \bar x= 65.6, \bar y= 30.9

Note that the least square estimates for a and b in Y= a+bx, are:

\hat{b}=\frac{\sum_{i=1}^n (x_i-\bar x)(y_i-\bar y)}{\sum_{i=1}^n (x_i-\bar x)^2}= \frac{-100.4}{290.4}=-0.3457

and \hat a= \bar{y}-\hat b\bar{x}=53.5779

a. What is the value of the slope of the regression line?

Ans. b= -0.3457

b. What is the value of the y-intercept?

Ans. a= 53.5779

c. Provide an interpretation of your answer to part b.

Ans. If X=0, that is, if the temperature would be 0 Fahrenheit, then the average number of employees who call in sick would be 53 approximately.

d. What is the final form of the regression line ?

Ans. Y= 53.5779 - 0.3457 x

e. Use Excel to obtain a scatterplot of the data with the regression line embedded in it.

Ans. Steps to obtain scatterplot with regression line embedded in it:

1) select columns of X and Y.

2) Go to insert -> charts -> scatter (X,Y)

3)Right click on data points, and select add trend line -> linear (you can also insert the trendline equation on the graph itself)

phpRWZi5v.png

f. What is the predicted number of employees that will call-in sick on a day that is 71 degrees Fahrenheit?

Ans. x=71,

then Y= 53.5779 - 0.3457 *71= 29 nearly.